OpenStudy (anonymous):

okay how do I do this? 1+(1/(1+(1/x))

OpenStudy (kc_kennylau):

continued to eternity or just 2 times?

OpenStudy (kc_kennylau):

\[1+\frac1{1+\frac1{1+\frac1{1+...}}}\] or \[1+\frac1{1+\frac1x}\]

OpenStudy (kc_kennylau):

oh wait sorry it terminated on the second term

OpenStudy (anonymous):

no just simplify it

OpenStudy (anonymous):

so yeah the 2nd one

OpenStudy (anonymous):

Whoops its 1 over x plus 1/x

OpenStudy (anonymous):

not 1 plus 1/x

OpenStudy (kc_kennylau):

\[1+\frac1{x+\frac1x}=1+\frac1{\frac{x^2+1}x}=1+\frac{x}{x^2+1}=\frac{x^2+x+1}{x^2+1}\]

OpenStudy (anonymous):

okay thanks a ton! I just wanted to see if I got it right ;)

OpenStudy (kc_kennylau):

no problem :)